I tried so many times to get the answer but i can't the problem is What percent of 16 is 5? The first thing i did i put 100 for the deominator then i put x for the numirator then i multiplied across. then i seen if it could be reduced and i still couldn't get my answer.
Answer by whisper(27)

There are two ways to solve this. Let's do the easy way first and use it as a check for the second.
First divide 5/16
.3125 or 31.25%
Next using a proportion:
x/100 = 5/16
Cross multiply:
16x = 5(100)
16x=500
500/16
31.25
31.25%
Please don't hesitate to ask if you have further questions!